Three consecutive even integers are such that the square of the first integer plus the square of the third integer is 136. Find the three integers.
Answer by checkley79(3056)

Let x, x+2 & x=4 be the three integers.
x^2+(x+4)^2=136
x^2+x^2+8x+16=136
2x^2+8x+16-136=0
2x^2+8x-120=0
2(x^2+4x-60)=0
2(x-6)(x+10)=0
x-6=0
x=6 ans. for the smallest integer.
6+4=10 ans. for the largest integer.
Proof:
6^2+10^2=136
36+100=136
136=136